# Joint Scheduling of DER under Demand Charges: Structure and Approximation

### v1

@@ Abstract
We study the joint scheduling of behind-the-meter distributed energy resources (DERs), including flexible loads, renewable generation, and battery energy storage systems, under net energy metering frameworks with demand charges. The problem is formulated as a stochastic dynamic program aimed at maximizing expected operational surplus while accounting for renewable generation uncertainty. We analytically characterize the structure of the optimal control policy and show that it admits a threshold-based form. However, due to the strong temporal coupling of the storage and demand charge constraints, the number of conditional branches in the policy scales combinatorially with the scheduling horizon, as it requires a look-ahead over future states. To overcome the high computational complexity in the general formulation, an efficient approximation algorithm is proposed, which searches for the peak demand under a mildly relaxed problem. We show that the algorithm scales linearly with the scheduling horizon. Extensive simulations using two open-source datasets validate the proposed algorithm and compare its performance against different DER control strategies, including a reinforcement learning-based one. Under varying storage and tariff parameters, the results show that the proposed algorithm outperforms various benchmarks in achieving a relatively small solution gap compared to the theoretical upper bound.
Index Terms:
Battery storage systems, demand charges, distributed energy resources, dynamic programming, energy management, flexible demands, net metering, peak searching.
@@ I Background and Literature
Utilities such as Pacific Gas and Electric (PG&E), Consolidated Edison (Con Edison), and Florida Power & Light (FPL) apply demand charges that vary based on customer class, time-of-use periods, and peak-coincident intervals [7, 8, 9]. These charges can constitute a substantial portion of monthly electricity bills, thereby incentivizing customers to strategically manage their load profiles. In response, utilities are adopting demand-based pricing not only to better reflect the cost of capacity provision but also to encourage peak load reduction and grid stability. From a system operator’s perspective, demand charges serve as a proxy for managing distribution-level constraints and capacity costs. Consequently, optimizing DER operations under such pricing schemes is essential for both economic efficiency and system reliability.
@@ I-A Related Work
Recent reinforcement learning (RL) advancements have further enriched the field. Lee et al. [20] introduced a novel approach to multi-residential energy scheduling using federated reinforcement learning, incorporating time-of-use and demand charge tariffs alongside energy storage and renewable sources. In parallel, the work in [21] developed a decision framework for demand-side management and NEM, aimed at peak load reduction and enhanced renewable energy utilization. Their work underscores the importance of managing flexible demands within the context of NEM.
@@
Despite these advancements, a gap remains in the literature for HEMS that co-optimize flexible demands and storage decisions under NEM and demand charges. Although [4, 11] addressed demand charges and [21] emphasized NEM and demand management, few studies combine all four elements. This work addresses this gap by proposing a HEMS that maximizes prosumer surplus under NEM, incorporating demand flexibility, storage scheduling, and demand charges, as summarized in Table I.
@@
This Work $\boldsymbol{\checkmark}$ $\boldsymbol{\checkmark}$ $\boldsymbol{\checkmark}$ $\boldsymbol{\checkmark}$
@@ I-B Summary of Results and Contributions
  • We show that the optimal policy maintains a threshold structure consistent with previous work [4] while demonstrating how the integration of multiple factors, such as demand flexibility, battery storage, and demand charges under bi-directional power flow leads to prohibitive computational costs of the optimal policy.
@@
The works in [4, 18] are most closely related to ours, as both adopt a stochastic dynamic programming framework for the prosumer decision problem. Our approach departs from [4] in two key aspects: (i) we model demand as flexible and jointly optimize it with battery storage, and (ii) we allow for bi-directional power and monetary flows between the prosumer and the DU under asymmetric import and export prices. Compared to [18], our model incorporates demand charges in addition to the volumetric charge, introducing stronger temporal coupling across decision stages. Furthermore, unlike both [4, 18], our work proposes an efficient approximation algorithm that achieves near-optimal performance, as validated by simulations against a theoretical upper bound.
I-C Mathematical Notations
[…]
@@ I-C Mathematical Notations
$\overline{\bm{d}}_{t}$ Consumption bundle’s upper limit.
@@ II Problem Formulation
Figure 1: DER setup with flexible demands $d\in\mathbb{R}_{+}$, renewable DG $g\in\mathbb{R}_{+}$, storage operation $e\in\mathbb{R}$, and net-consumption $z\in\mathbb{R}$. The net energy metering tariff with demand charges is also demonstrated with cost associated with amount of net consuming, net producing and the peak load over the period of day where peak load is recorded.
We study a HEMS for prosumers that co-optimizes flexible demand, storage operation, and renewable distributed generation under a NEM tariff with distinct buy/sell rates and peak demand charges. The optimization objective maximizes the operational surplus defined as consumption utility minus electricity costs, where demand charges are levied on the maximum consumption drawn from either load, battery charging, or both. As shown in Fig. 1, the energy storage and flexible loads are jointly scheduled to reduce the demand charge by shaving the load curve, and reduce the volumetric charge by arbitraging NEM rates, while still satisfying prosumer’s consumption.
@@ II-A Prosumer Resources
The flexible demands are denoted by the vector
@@
where $K$ is the total number of prosumer devices. Here $\overline{\bm{d}}$ is the consumption bundle’s upper limit.
[…]
The prosumer also owns a BTM battery energy storage, which has output $e_{t}\in[-\underline{e},\overline{e}]$, where $\underline{e}$ and $\overline{e}$ are discharging and charging limits. The output is decomposed into charging $[e_{t}]^{+}=\max\{e_{t},0\}$ and discharging $[e_{t}]^{-}=\max\{-e_{t},0\}$. With charge/discharge efficiencies $\tau,\rho\in(0,1]$, the state of charge (SoC) $s_{t}\in[0,B]$ evolves as:
@@ II-B System Dynamics
The system state at time $t$, $x_{t}$ is described by $x_{t}=[s_{t},g_{t},c_{t}]$, where $c_{t}$ is the prosumer’s peak demand before time $t$. We allow the system to be controlled by battery charging/discharging and consumption adjustments, i.e., the control action is $u_{t}\coloneqq(e_{t},\bm{d}_{t})$. The peak demand in the state evolves as
@@
II-C NEM X Tariff with Demand Charges
Under NEM, when the prosumer’s gross loads (from consumption and storage charging) are higher than its generation (from renewable DG), the prosumer pays at the retail rate. Otherwise, the prosumer gets compensated at the export rate.
Let $(p_{t}^{+}\geq p_{t}^{-},p)\geq 0$ be the retail, export, and peak demand prices, respectively. We equivalently break down the peak demand charge at the end of time $T$ to each time step $t$ since $pc_{T}=p\sum_{t=0}^{T-1}[z_{t}-c_{t}]^{+}$. Thus, we can write the payment (cost) under NEM X in [3] for each time $t$ as
$P_{t}(z_{t},c_{t})=p_{t}^{+}[z_{t}]^{+}-p_{t}^{-}[z_{t}]^{-}+p[z_{t}-c_{t}]^{+},$ (5)
where $P_{t}(\cdot)>0$ ($P_{t}(\cdot)<0$) indicates that the prosumer pays (gets compensated) to (by) the DU.
@@ II-E DER Co-Optimization Problem
To optimally schedule its DER, the prosumer solves a dynamic program that maximizes its reward function by explicitly accounting for the temporal coupling of storage operations and an NEM tariff with bidirectional energy pricing and peak demand charges.
@@
$r_{t}(x_{t},u_{t})=r_{t}((s_{t},g_{t},c_{t}),(e_{t},\mathbf{d}_{t}))=U_{t}(\bm{d}_{t})-P_{t}(z_{t},c_{t}),\quad t=0,\ldots,T-1.$ (7)
@@
The policy $\pi=(\pi_{0},\ldots,\pi_{T-1})$ is a sequence of mapping from state space to action space such that $\pi_{t}(x_{t})=u_{t}$. Given initial state $x_{0}=(s,g,0)$, the storage-consumption co-optimization is then defined as:
@@
Subject to for $\quad\text{for all}\ t=0,1,\ldots,T-1$
@@
Constraints (9e) and (9h) induce intertemporal coupling, precluding a myopic (stage-wise optimal) policy. Specifically, relaxing either constraint in isolation is insufficient to eliminate the temporal dependencies, and the resulting optimal policy remains non-myopic (see Thm. 4 in the Appendix). However, when both constraints are simultaneously relaxed, the problem admits a myopic optimal policy characterized by decoupled per-period decisions (see Lem. 4).
@@ III Dynamic Program Analytical Results
[…]
The following lemma establishes a conclusion regarding the concavity of the optimal value function.
@@
[…]
First, by the concavity of $U_{t}$:By definition, $U_{t}$ is concave:
@@
[…]
The cost $P_{t}(z_{t},c_{t})$ consists of two components: the energy cost $p_{t}^{+}[z_{t}]^{+}-p_{t}^{-}[z_{t}]^{-}$ and the demand charging $p[z_{t}-c_{t}]^{+}$. Since $p_{t}^{+}\geq p_{t}^{-}\geq 0$, the energy cost is convex.We separate the $P_{t}(z_{t},c_{t})$ into energy cost $p_{t}^{+}[z_{t}]^{+}-p_{t}^{-}[z_{t}]^{-}$ and the demand charging cost $p[z_{t}-c_{t}]^{+}$. By assumption, $p_{t}^{+}\geq p_{t}^{-}\geq 0$, and $p_{t}^{+}[0]^{+}=p_{t}^{-}[0]^{-}=0$, the left derivative at 00 is smaller than the right derivative at 00. So, the derivative of $p_{t}^{+}[z_{t}]^{+}-p_{t}^{-}[z_{t}]^{-}$ is non-decreasing and hence $p_{t}^{+}[z_{t}]^{+}-p_{t}^{-}[z_{t}]^{-}$ is convex. Then,
$-\sum_{t=\tau}^{T}(p_{t}^{+}[z_{t}]^{+}-p_{t}^{-}[z_{t}]^{-}+p_{t}^{+}[z_{t}^{\prime}]^{+}-p_{t}^{-}[z_{t}^{\prime}]^{-})/2$ $\leq-\sum_{t=\tau}^{T}p_{t}^{+}[(z_{t}+z_{t}^{\prime})/2]^{+}-p_{t}^{-}[(z_{t}+z_{t}^{\prime})/2]^{-}$ (11)
@@
$-\sum_{t=\tau}^{T}(p[z_{t}-c_{t}]^{+}+p[z_{t}^{\prime}-c_{t}^{\prime}]^{+})/2=-p(\max_{\tau\leq t\leq T}\{z_{t}\}-c_{\tau}+\max_{\tau\leq t\leq T}\{z_{t}^{\prime}\}-c_{\tau}^{\prime})/2$ $\leq-p(\max_{\tau\leq t\leq T}\{(z_{t}+z_{t}^{\prime})/2\}-(c_{\tau}+c_{\tau}^{\prime})/2)$ (12)
@@
[…]
@@
Building on the results of Lem. 1, Lem. 2, and Lem. 3, we establish Thm. 1, which characterizes the structure of the optimal co-optimization policy under demand charges.
@@
[…]
@@
[…]
must exist in the interior of the feasible set $\mathcal{F}_{t}$ or at its boundary. The optimal action can be obtained by solving concave optimizations with and without boundary constraints, which results in a threshold policy. So, we can get the optimal solution by solving $(K+1)2^{K}$ constrained concave optimization and one unconstrained concave optimization, which results in a threshold solution of $(K+1)2^{K}+1$ cases. ∎
Although Thm. 1 reveals a structured policy, direct dynamic programming implementation remains impractical. First, optimization necessitates gradients of $r_{t}$ and $Q_{t}$, which are computationally prohibitive for complex utility functions. Also, continuous-space Q-function tracking requires costly discretization, as validated in simpler settings [4]. To this end, we develop an efficient approximation algorithm to overcome these computational barriers.
IV Efficient Approximation Algorithm
Since computing the exact optimal solution for problem (9) is computationally intractable, we propose an efficient approximation algorithm. The large storage peak searching (LSPS) algorithm efficiently searches for the optimal peak demand by solving a relaxed version of the original problem. The resulting solution is then projected onto the feasible region $\mathcal{F}_{t}$, producing near-optimal and computationally tractable approximations. The original formulation prevents myopic solutions through constraints (9e) and (9h). To address this computational challenge, we first relax constraint (9e) by considering a large capacity regime (Sec. IV-A), then develop an efficient method to handle constraint (9h) (Sec. IV-B).

### v2

@@ Abstract
We study the joint scheduling of behind-the-meter distributed energy resources (DERs), including flexible loads, renewable generation, and battery energy storage systems, under net energy metering tariffs with demand charges. The problem is formulated as a stochastic dynamic program aimed at maximizing expected operational surplus while accounting for renewable generation uncertainty. We analytically characterize the optimal control policy and show that it admits a threshold-based structure. However, due to the strong temporal coupling of the storage and demand charge constraints, the number of conditional branches in the policy scales combinatorially with the scheduling horizon, as it requires a look-ahead over future states. To overcome the high computational complexity in the general formulation, an efficient approximation algorithm is proposed, which searches for the peak demand under a mildly relaxed problem. We show that the algorithm scales linearly with the scheduling horizon. Extensive simulations using two open-source datasets validate the proposed algorithm and compare its performance against different DER control strategies, including a reinforcement learning-based one. Under varying storage and tariff parameters, the results show that the proposed algorithm outperforms various benchmarks in achieving a relatively small solution gap compared to a theoretical upper bound.
@@ I Background and Literature
Utilities such as Pacific Gas and Electric (PG&E), Consolidated Edison (Con Edison), Florida Power & Light (FPL), and Arizona Public Service (APS) apply demand charges that vary based on customer class, time-of-use periods, and peak-coincident intervals [7, 8, 9, 10]. These charges can constitute a substantial portion of monthly electricity bills, thereby incentivizing customers to strategically manage their load profiles. In response, utilities are adopting demand-based pricing not only to better reflect the cost of capacity provision but also to encourage peak load reduction and grid stability. From a system operator’s perspective, demand charges serve as a proxy for managing distribution-level constraints and capacity costs. Consequently, optimizing DER operations under such pricing schemes is essential for both economic efficiency and system reliability.
@@ I-A Related Work
Recent progress in reinforcement learning (RL) has further enriched the area by enabling data-driven approaches to jointly optimize DER under complex tariffs. For example, Lee et al. [22] introduced a federated RL framework for multi-residential energy scheduling that incorporates time-of-use and demand charge tariffs alongside energy storage and renewable sources. Similarly, [23] developed a decision framework for demand-side management under NEM, aiming at peak load reduction and improved renewable utilization. While RL offers scalability and adaptability to uncertain environments, it generally lacks the interpretability and optimality guarantees of dynamic programming, which can make it difficult to validate solutions or ensure consistent performance across varying operating conditions.
Despite these advancements, a gap remains in the literature for HEMS that co-optimize flexible demands and storage decisions under NEM and demand charges. Although [4, 12] addressed demand charges and [23] emphasized NEM and demand management, few studies combine all four elements. This work addresses this gap by proposing a HEMS that maximizes prosumer surplus under NEM, incorporating demand flexibility, storage scheduling, and demand charges, as summarized in Table I, and explained in the next section.
@@
[…]
@@ I-B Summary of Results and Contributions
  • We show that the optimal policy maintains a threshold structure consistent with previous work [4] while demonstrating how the integration of multiple factors, such as demand flexibility, battery storage, and demand charges under bi-directional power flow, leads to prohibitive computational costs of the optimal policy.
@@
The works in [4, 19] are most closely related to ours, as both adopt a stochastic dynamic programming framework for the prosumer decision problem. Our approach departs from [4] in two key aspects: (i) we model demand as flexible and jointly optimize it with battery storage, and (ii) we allow for bi-directional power and monetary flows between the prosumer and the DU under asymmetric import and export prices. Symmetric NEM import and export prices, i.e., NEM 1.0, are known to create utility revenue shortfalls [24] potentially leading to death spirals [25] and cross-subsidies [26]. Consequently, several states, including California, Arizona, Illinois, Georgia, and Utah, have adopted successor NEM tariffs featuring asymmetric import and export prices [27].
Compared to [19], our model incorporates demand charges in addition to the volumetric charge, introducing stronger temporal coupling across decision stages. From an economic perspective, demand charges improve the cost reflectivity of the tariff and reduce the over-reliance on volumetric charges to recover bundled cost components [24].
Lastly, unlike both [4, 19], our work proposes an efficient approximation algorithm that outperforms various benchmarks in achieving a relatively small solution gap compared to a theoretical upper bound.
I-C Mathematical Notations and Paper Organization
Table II lists the major symbols used. Throughout the paper, boldface letters indicate column vectors as in $\bm{x}=(x_{1},\ldots,x_{n})$ with $\bm{x}^{\top}$ as its transpose. In particular, $\bm{1}$ is a column vector of all ones. For a multivariate differentiable function $f$ of $\bm{x}$, we use interchangeably $f(\bm{x})$ and $f(x_{1},\ldots,x_{n})$, and we use $f^{\prime}$ and $f^{\prime\prime}$ to denote its first-order and second-order derivative (if it exists), respectively. We use $f^{-1}$ to denote its inverse function (if it exists). For vectors $\bm{x},\bm{y}$, $\bm{x}\preceq\bm{y}$ is the element-wise inequality $x_{i}\leq y_{i}$ for all $i$, and $\bm{x}]^{+},[\bm{x}]^{-}$ are the element-wise positive and negative parts of vector $\bm{x}$, i.e., $x_{i}]^{+}:=\max\{0,x_{i}\}$, $x_{i}]^{-}:=-\min\{0,x_{i}\}$ for all $i$, and $\bm{x}=[\bm{x}]^{+}-[\bm{x}]^{-}$. We denote by $\mathbb{R}_{+}:=\{x\in\mathbb{R}:x\geq 0\}$ the set of non-negative real numbers.
The rest of the paper is organized as follows. §II, provides a mathematical formulation of DERs, tariff under demand charge, system dynamics, and the joint scheduling decision problem. §III presents some analytical results on the dynamic programming problem, and §IV proposes an efficient approximation algorithm that circumvents the exponential complexity of solving the dynamic program. A detailed case study using real data from two buildings is presented in §V, followed by concluding remarks and future directions in §VI.
@@ I-C Mathematical Notations and Paper Organization
$\overline{\bm{d}}$ Consumption bundle’s upper limit.
@@ II Problem Formulation
Figure 1: DER setup with flexible demand $d\in\mathbb{R}+$, renewable generation $g\in\mathbb{R}+$, storage operation $e\in\mathbb{R}$, and net consumption $z\in\mathbb{R}$. The illustrated NEM tariff includes both energy and demand charges, with costs associated with net consumption, net generation, and the peak load recorded over the daily billing period.
We study a HEMS for prosumers that co-optimizes flexible demand, storage operation, and renewable distributed generation under a NEM tariff with distinct buy/sell rates and peak demand charges. The optimization objective maximizes the operational surplus defined as consumption utility minus electricity costs, where demand charges are levied on the maximum consumption drawn from either load, battery charging, or both. As shown in Fig. 1, the energy storage and flexible loads are jointly scheduled to reduce the demand charge by shaving the load curve, and reduce the volumetric charge by arbitraging NEM rates, while still satisfying the prosumer’s consumption.
@@ II-A Prosumer Resources
The prosumer flexible demands are denoted by the vector
@@
where $K$ is the total number of devices, and $\overline{\bm{d}}$ is the consumption bundle’s upper limit. [^3] The minimum consumption limit is set to $\bm{0}$ without loss of generality. Some important devices might be less flexible or inflexible. Mathematically, this is incorporated by assigning these devices stricter flexibility limits, i.e., for an inflexible device $k$, whose consumption should not deviate from $\tilde{d}_{t}^{k}$, its flexibility limits would be $\tilde{d}_{t}^{k}\leq d_{t}^{k}\leq\tilde{d}_{t}^{k}$.
The prosumer also owns a BESS, which has output $e_{t}\in[-\underline{e},\overline{e}]$, where $\underline{e}$ and $\overline{e}$ are the discharging and charging limits, respectively. The output is composed of charging $e_{t}]^{+}=\max\{e_{t},0\}$ and discharging $e_{t}]^{-}=\max\{-e_{t},0\}$ actions. Given charging and discharging efficiencies, $\tau\in(0,1]$ and $\rho\in(0,1]$, respectively, the BESS state of charge (SoC) $s_{t}\in[0,B]$ evolves as:
@@ II-B System Dynamics
The system state at time $t$, $x_{t}$, is described by $x_{t}=[s_{t},g_{t},c_{t}]$, where $c_{t}$ is the prosumer’s peak demand before time $t$. The system is controlled by the BESS action, i.e., charging or discharging, and consumption adjustments, hence the control action is $u_{t}\coloneqq(e_{t},\bm{d}_{t})$. The peak demand in the state evolves as
@@
II-C NEM Tariff with Demand Charges
Under NEM, when the prosumer’s gross load (from consumption and storage charging) is higher than its generation (from renewable DG and storage discharging), the prosumer pays at the retail rate. Otherwise, the prosumer gets compensated at a predetermined export rate. To better reflect the capacity costs incurred by the DU, a demand charge is added to the prosumer payment. The demand charge is a one-time payment based on the peak demand over the billing period.
Let $(p_{t}^{+}\geq p_{t}^{-},p)\geq 0$ be the retail, export, and peak demand prices, respectively. The payment under NEM $P_{t}(\cdot)$ over the considered time horizon is given by
$\sum_{t=0}^{T-1}P_{t}(z_{t},c_{t})=\sum_{t=0}^{T-1}\left(p_{t}^{+}[z_{t}]^{+}-p_{t}^{-}[z_{t}]^{-}\right)+pc_{T}+A,$ (5)
where $A$ denotes a fixed charge, commonly known as a connection charge, that does not vary with consumption [3]. The payment in (5) corresponds to a three-part tariff structure, consisting of volumetric, demand, and fixed components [28].
For the convenience of analysis, we equivalently break down the peak demand charge at the end of time $T$ to each time step $t$ as a rolling-based incremental charge, since
[…]
Thus, we can write the payment under NEM X in (5) for each time $t$ as
$P_{t}(z_{t},c_{t})=p_{t}^{+}[z_{t}]^{+}-p_{t}^{-}[z_{t}]^{-}+p[z_{t}-c_{t}]^{+}+A_{t},$ (6)
where, for every time $t$, $P_{t}(\cdot)>0$ ($P_{t}(\cdot)<0$) indicates that the prosumer pays (gets compensated) to (by) the DU. [^4] We assume $A_{t}=0,\forall t$, since fixed costs do not influence DER scheduling decisions, at least in the short-run.
Special cases of the general NEM tariff in (6) are considered in [19] when demand charges ($p[z_{t}-c_{t}]^{+}$) are absent and in [4] when energy exports are absent ($p_{t}^{-}[z_{t}]^{-}$).
@@ II-E DER Co-Optimization Problem
To optimally schedule its DER, the prosumer solves a dynamic program that maximizes its reward function by explicitly accounting for the temporal coupling of storage operations and a NEM tariff with asymmetric retail and export rates and peak demand charges.
@@
$r_{t}(x_{t},u_{t})$ $=r_{t}((s_{t},g_{t},c_{t}),(e_{t},\mathbf{d}_{t}))$ (8)
$=U_{t}(\bm{d}_{t})-P_{t}(z_{t},c_{t}),\quad t=0,\ldots,T-1.$
@@
The policy $\bm{\pi}=(\pi_{0},\ldots,\pi_{T-1})$ is a sequence of mappings from state space to action space such that $\pi_{t}(x_{t})=u_{t}$, for every $t$. Given initial state $x_{0}=(s,g,0)$, the storage-consumption co-optimization is then defined as:
@@
Subject to $\quad\text{for all}\ t=0,1,\ldots,T-1$
@@
Constraints (10e) and (10h) induce intertemporal coupling, precluding a myopic (stage-wise) optimal policy. Specifically, relaxing either constraint in isolation is insufficient to eliminate the temporal dependencies, and the resulting optimal policy remains non-myopic (see Theorem 4 in Appendix -A). However, when both constraints are simultaneously relaxed, the problem admits a myopic optimal policy characterized by decoupled per-period decisions (see Lemma 4).
@@ III Dynamic Program Analytical Results
In this section, we show that the optimal policy for the co-optimization problem in §II-E is a threshold-based one.
First, we show that the optimal value function is concave.
@@
Consider two different states $x_{\tau}$ and $x_{\tau}^{\prime}$ with their corresponding optimal trajectories $\{(x_{t},u_{t})\}_{t=\tau}^{T}$ and $\{(x_{t}^{\prime},u_{t}^{\prime})\}_{t=\tau}^{T}$. Let $\tilde{x}_{\tau}=(x_{\tau}+x_{\tau}^{\prime})/2$ be their average initial state, and define the average policy $\tilde{\pi}=\{\tilde{u}_{t}=(u_{t}+u_{t}^{\prime})/2\}_{t=\tau}^{T}$ with corresponding trajectory $\{\tilde{x}_{t}\}_{t=\tau}^{T}$.
First, by the concavity of $U_{t}$:
@@
The cost $P_{t}(z_{t},c_{t})$ consists of two components: the energy cost $p_{t}^{+}[z_{t}]^{+}-p_{t}^{-}[z_{t}]^{-}$ and the demand charging $p[z_{t}-c_{t}]^{+}$. Since $p_{t}^{+}\geq p_{t}^{-}\geq 0$, the energy cost is convex. Then,
$-\sum_{t=\tau}^{T}(p_{t}^{+}[z_{t}]^{+}-p_{t}^{-}[z_{t}]^{-}+p_{t}^{+}[z_{t}^{\prime}]^{+}-p_{t}^{-}[z_{t}^{\prime}]^{-})/2$ (12)
$\leq-\sum_{t=\tau}^{T}p_{t}^{+}[(z_{t}+z_{t}^{\prime})/2]^{+}-p_{t}^{-}[(z_{t}+z_{t}^{\prime})/2]^{-}$
@@
$-\sum_{t=\tau}^{T}(p[z_{t}-c_{t}]^{+}+p[z_{t}^{\prime}-c_{t}^{\prime}]^{+})/2$ (13)
$=$ $-p(\max_{\tau\leq t\leq T}\{z_{t}\}-c_{\tau}+\max_{\tau\leq t\leq T}\{z_{t}^{\prime}\}-c_{\tau}^{\prime})/2$
$\leq$ $-p(\max_{\tau\leq t\leq T}\{(z_{t}+z_{t}^{\prime})/2\}-(c_{\tau}+c_{\tau}^{\prime})/2)$
@@
The optimal value function $V_{t}^{*}(s_{t},g_{t},c_{t})$ is non-decreasing in $s_{t}$, $g_{t}$, and $c_{t}$ for every time step $t\leq T$.
@@
Building on the results of Lemma 1, Lemma 2, and Lemma 3, we establish Theorem 1, which describes the structure of the optimal co-optimization policy under demand charges.
@@
The feasible region for action $u_{t}$, $(e_{t},\bm{d}_{t})\in\mathcal{F}_{t}=[\max\{-\underline{e},-\rho s_{t}\},\min\{\overline{e},(B-s_{t})/\tau\}]\times\prod_{i=1}^{K}[0,d_{i}]$ is a compact set. From Lemma 1 to Lemma 3, the optimal value function and $Q$-function is concave in action $u_{t}$, so the optimal action
@@
must exist in the interior of the feasible set $\mathcal{F}_{t}$ or at its boundary. The optimal action can be obtained by solving concave optimizations with and without boundary constraints, which results in a threshold policy. ∎
Although Theorem 1 reveals a structured policy, direct dynamic programming implementation remains impractical. If we directly apply concave optimization on the compact sets, it necessitates gradients of $r_{t}$ and $Q_{t}$, which are computationally prohibitive, especially for complex utility functions that may not be additive with respect to consuming devices. On the other hand, if we want to avoid the expensive optimization by backtracking, the continuous-space $Q$-function tracking requires costly discretization, as validated in simpler settings in [4]. To this end, we develop, in §IV, an efficient approximation algorithm to overcome these computational barriers.
IV An Efficient Approximation Algorithm
Since obtaining the exact optimal solution to problem (10) is computationally intractable, we develop an efficient approximation algorithm. The proposed large storage peak searching (LSPS) algorithm efficiently searches for the optimal solution by solving a relaxed version of the original problem. The resulting solution is then projected onto the feasible region $\mathcal{F}_{t}$, producing near-optimal and computationally tractable approximations. The original formulation prevents myopic solutions through constraints (10e) and (10h). To address this computational challenge, we first relax constraint (10e) by considering a large storage capacity regime (see §IV-A), then develop an efficient method to handle constraint (10h) (see §IV-B).